OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = eight divided by x squared + 2 HELP PLEASE

OpenStudy (anonymous):

you have lots of choices you could say \(g(x)=x^2, f(x)=\frac{8}{x}\) and then \[f(g(x))=f(x^2)=\frac{8}{x^2}\] but that is just one option

OpenStudy (anonymous):

Ok thank you so much!

OpenStudy (anonymous):

yw